OpenStudy (anonymous):

1. At a fundraiser breakfast, the bill for three glasses of orange juice and five pancake specials is $7.60, whereas the bill for one glass of orange juice and two pancake specials is $2.90. a. Write a system of two equations to model this problem. Let j stand for the price of a glass of juice and let p stand for the price of a pancake special. b. Use the substitution method to solve the system of equations. c. What would be the bill for one glass of orange juice and one pancake special?

OpenStudy (anonymous):

let price of juice= j let price of pancake=p => 3j+5p=7.60 j+2p=2.90

OpenStudy (anonymous):

from 2nd equation we have j=2.90-2p substituting j into equation 1 we have 3(2.90-2p)+5p=7.60 solve this for the value of p and put it in the equation(j=2.90-2p) to find the value of j this will give u the price of juice and pancakes to solve the third part... :)

OpenStudy (anonymous):

can u solve it for me?

OpenStudy (anonymous):

ok w8 sorry OS is laggin much